# Leveraging Diffusion Models for Parameterized Quantum Circuit Generation

### v2

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
The promise of quantum computing depends on the efficient design of quantum circuits that can be executed on currently available hardware. In the present noisy intermediate-scale quantum (NISQ) [2] era, where this challenge is most acute, circuit synthesis must account for hardware constraints such as limited qubit counts, restricted gate sets, and connectivity, while still achieving target performance metrics [3, 4, 5]. The design of an appropriate ansatz for a given task – be it approximating a target unitary, solving an optimization problem or learning from data for a quantum machine learning (QML) task – is challenging and typically relies on human expertise or brute-force search. In recent years, there has been growing interest in leveraging heuristic optimization and machine learning (ML) techniques to automate quantum circuit design. Prior approaches include evolutionary strategies [6, 7, 8, 9], reinforcement learning [10, 11, 4, 12], Bayesian optimization [13], and adaptive algorithms [14, 15]. While these methods can discover novel circuit designs, they typically require repeated evaluations of quantum metrics or cost functions during training [16][17]. This evaluation process becomes intractable for large systems and trainable parameterized quantum circuits (PQCs), creating a significant computational bottleneck for automated circuit synthesis methods [18]. Moreover, techniques trained on specific quantum tasks or circuit layouts often show poor generalization to new scenarios, as they tend to encode task-specific features rather than learning transferable principles [19]. To address these challenges, [1] introduce the idea of applying denoising diffusion models (DM) to quantum circuit synthesis. The proposed text-conditioned diffusion pipeline generates gate-based circuits for a given target objective without explicitly simulating quantum dynamics during training. While this work established DMs as a promising tool for quantum circuit synthesis, an important open challenge is handling continuous gate parameters. Commonly used rotation gates, such as single-qubit or controlled rotations, are parameterized by continuous values and significantly impact the functionality of PQCs.
In this paper, we introduce an extended DM-based pipeline that integrates continuous-valued gate parameters alongside discrete gate choices into the generation process. We train the proposed model architecture on extensive datasets of quantum circuits with varying gate sets, designed for state preparation and QML classification tasks. Experimental results demonstrate that the described methodology effectively bypasses the computationally expensive parameter training phase during the search, addressing a key limitation of many existing approaches to the the automated design of PQCs, and shows strong generalization ability to new scenarios. At inference time, the model driven by a textual prompt produces PQCs that, when subject to specified constraints, are immediately deployable on the target quantum device.
@@ II Related Work
The approach introduced in this work extends [1], which treat quantum circuits as structured data samples and train a generative denoising diffusion model (DM) to design new circuit designs. DMs generate synthetic data by iteratively denoising random noise samples through a learned reverse diffusion process. Introduced by [20], the fundamental principle involves adding Gaussian noise to training samples as forward Markovian diffusion process, then training a neural network to reverse noising procedure and approximate the data distribution’s score function. After training, data generation begins from a pure noise sample, and the trained model iteratively predicts and removes noise, step by step, resulting in samples resembling those from the training dataset. DMs employ denoising autoencoder architectures implemented through neural networks, typically structured to predict either the original noise-free data or the noise itself. Recent advances, such as latent diffusion and classifier-free guidance, have made DMs a leading approach in conditional generative modeling tasks across numerous domains, including image and video generation [21, 22], audio synthesis [23], and protein structure prediction [24]. Extending these advances, [1] explored the potential of DMs beyond classical domains and demonstrated their effectiveness in the context of quantum circuit synthesis. Specifically, the authors applied a denoising DM to two quantum computing tasks: (1) entanglement generation, where the model produced circuits that realize specific multi-qubit entangled states, and (2) unitary compilation, where the model generated circuits approximating random target unitaries using constrained gate sets. The DM excelled in generating valid, novel circuits and enabled advanced capabilities such as masking and editing parts of a circuit post hoc without retraining. Notably, the approach eliminates the need for classical simulations to compute cost functions comparing generated circuits to target unitaries. Ref. [25] proposes an architectural enhancement by replacing the standard U-Net used for noise prediction with a U-Net-style Diffusion Transformer. A recent concurrent work [26] applied a diffusion model to generate ansätze for variational quantum eigensolver (VQE), with a focus on quantum chemistry problems. The model was trained on a corpus of unitary coupled cluster (UCC) ansatz circuits encoded into images and successfully proposed new circuits that achieved comparable ground-state energies in VQE simulations. However, the techniques proposed in earlier studies represent gates using fixed-dimensional vectors drawn from a finite embedding set [27], limiting its ability to express parametrized quantum operations. In this work, we extend the methodology to enable the generation of PQCs with continuous gate parameters.
@@
Beyond diffusion-based models, this work also relates more broadly to the field of quantum architecture search (QAS) research, where the goal is to automatically generate PQCs tailored to specific tasks and hardware constraints [28]. While the most approaches proposed for QAS treat the design of PQCs as a two-step process [19, 4, 17, 13], i.e., first selecting discrete gate sequences and then using classical optimizers to determine continuous parameters, several recent methods aim to learn both circuit structures and parameterizations simultaneously, optimizing over both discrete gates and continuous parameters in a unified pipeline. Ref. [29] incorporates the parameter optimization task into reinforcement learning settings enabling continuous actions. The generative quantum eigensolver (GQE) proposed in [30] is a transformer-based approach for generating of PQCs. This approach is designed for chemical problem of searching for the ground state of the electronic structure Hamiltonians of several molecules. Similarly to our approach, it integrates continuous optimizable gate parameters into a classical generative model. However, unlike the transformer-based model in GQE, we employ a denoising diffusion model for PQCs generation, which iteratively refines random noise samples to generate circuits, offering advantages in generating diverse quantum circuits with novel structures.
@@ III-A Circuit Encoding
Figure 1: Overview of our diffusion-based PQC synthesis pipeline, inspired by [1] (a) Circuit Encoding: A quantum circuit is converted into a tensor by concatenating the embeddings of discrete gates and continuous parameters (e.g., rotation angles). (b) Forward Diffusion: Gaussian noise is iteratively added to this tensor, simulating a corruption process. (c) Textual Conditioning: A pretrained CLIP encoder converts task-specific prompts into embeddings. (d) Reverse Diffusion: A U-Net with cross-attention receives the noisy tensor and text embedding, iteratively removing noise to restore gate identities and parameters. (e) Inference and Decoding: At inference, random noise is denoised under the given prompt, and the final tensor is mapped back into a readable quantum circuit.
To train our diffusion model, we represent each PQCs using two structured tensors, as depicted in Fig. 1. The first tensor is three-dimensional, where one dimension corresponds to qubit indices $N$, another to discrete time steps $T$, and the third dimension encodes discrete gate identities $d_{c}$. For each discrete times $t_{i}\in T$ step, only one gate per qubit is allowed. If no gate acts on a qubit at a particular timestep (i.e., an idle operation), we assign a special no-op token. To convert this discrete representation into a continuous embedding suitable for the diffusion model, each discrete gate type is transformed into a fixed-length embedding vector determined before training. This embedding process results in a continuous tensor representation of gates across qubits and time steps. To handle gates with continuous parameters, e.g., a rotation gate $rx(\theta)$, we introduce a second tensor with dimensions corresponding to qubit indices, discrete time steps, and the maximal number of normalized parameters used by any gate in the circuit $d_{p}$. In this parameter tensor, we encode the numerical values of gate parameters, placing a no-op token for gates without parameters. By concatenating these two tensors, each gate operation at every qubit and timestep is uniquely encoded by its discrete type embedding and its continuous parameter. As a result, identical gates with identical parameters map to identical points in this combined embedding space, whereas gates differing in type or parameter values map distinctly. This method naturally extends to gates with multiple parameters. During decoding, the combined tensor is split into a circuit matrix and a parameter matrix. For each qubit–timestep position in the circuit matrix, we use cosine similarity to match the generated gate embedding to its closest counterpart among the known gate embeddings. The corresponding parameter matrix is then applied only when a gate requires parameters, using the remaining $d_{p}$ dimensions of the vector as a parameter encoding. In principle, if a gate supports multiple parameters (not present in our current gate sets), one could introduce an inverse mapping or train an auxiliary decoder network to handle this more complex parameter space.. If the chosen gate $G$ is non-parameterized (e.g., an ’h’ gate), the parameter subvector is simply ignored.
@@ III-B Model Architecture
In this work, we extend the quantum circuit synthesis pipeline described by [1]. Specifically, we introduce textual conditioning via a pre-trained CLIP text encoder - originally trained to align visual and textual representations through Contrastive Language-Image Pre-training (CLIP) - to guide our diffusion model towards generating PQCs optimized for specific tasks, i.e., quantum state fidelity or classification accuracy maximization. Task-specific textual prompts, for instance "Generate GHZ fidelity: $\mathsf{fid\_value}$” or "Generate Accuracy: $\mathsf{acc\_value}$”, are encoded into continuous embeddings to condition the diffusion process.
The core of our diffusion model utilizes a U-Net architecture, widely adopted in state-of-the-art diffusion-based image synthesis. The model accepts as input a noisy quantum circuit tensor of dimensions $(d_{c}+d_{p})\times N\times T$, representing embedding dimensions for gates and parameters, qubits and time steps, respectively. It outputs a noise-residual tensor of matching shape. To enable textual conditioning, we augment the standard U-Net structure with cross-attention layers, integrating CLIP-generated textual embeddings. This design explicitly conditions the generated circuits on task-specific textual prompts, eliminating the need for explicit labels of fidelity or accuracy, which the model implicitly learns from training data alone. Note that currently our approach supports parameter embeddings with $d_{p}\leq 1$.
@@ III-C Training Strategy
Our DM is trained by simulating a forward diffusion process on quantum circuit data and learning its reverse denoising counterpart in a manner consistent with DDPM  [31] procedures. Each training iteration involves selecting a quantum circuit encoding $x_{0}$, adding Gaussian noise at a randomly sampled diffusion timestep $t$ to obtain a noised version $x_{t}$, and training the U-Net to predict this added noise, $\hat{\epsilon}_{\theta}(x_{t},t,\texttt{prompt})$. Model parameters are optimized by minimizing the mean-squared error $\mathcal{L}~{}\;=~{}\;\mathbb{E}_{x_{0},\epsilon,t}\bigl{[}\|\epsilon~{}\;-~{}\hat{\epsilon}_{\theta}(x_{t},t,\texttt{prompt})\|^{2}\bigr{]}.$
@@
Conditioning on textual prompts (e.g., ‘”Generate GHZ fidelity: 1.0000”‘) is done through cross-attention layers, where the prompt is embedded by a CLIP encoder (output size $77\times 512$). For numerical stability, fidelity prompts are rounded to four decimal places, while ML accuracy prompts use one decimal place. During training, with a 10% probability, the prompt is replaced by an empty token. This “unconditional” path enables  classifier-free guidance (CFG) at inference, whereby we can smoothly interpolate between conditional and unconditional predictions  [32]. We train for $T=1000$ diffusion steps using Adam with a one-cycle learning-rate schedule.
At inference time, we draw Gaussian noise as an initial circuit encoding and iteratively denoise it using the trained U-Net. When classifier-free guidance is enabled, we selectively mix conditional and unconditional predictions to control how strictly the model follows the prompt. The final denoised tensor is decoded into a human-readable quantum circuit: we pick the closest discrete gate embedding at each position and apply the predicted continuous parameters (if any) to obtain the final PQC.
@@ IV-A Dataset
A key element of our approach is the construction of large training datasets consisting of quantum circuits, which the diffusion model leverages to learn the circuit performance distribution. For two tasks we address, i.e., quantum state preparation and classification accuracy optimization, we build on the existing SQuASH [^1] Availibale at https://github.com/SQuASH-bench/SQuASH dataset [33]. Specifically, for the quantum state preparation, the dataset includes quantum circuits designed to generate a $3$-qubit GHZ state within a predefined approximation error. The data were collected from various search strategies used to generate circuits for the task, including random search, reinforcement learning, and an evolutionary approach. To demonstrate the model’s adaptability to different hardware constraints, we define two distinct data subsets, each characterized by a different gate alphabet. The first subset incorporates PQCs constructed using the gate set $gs1$ = {’cx’, ’h’, ’rx’, ’ry’, ’rz’, ’id’}. The second subset uses  $gs2$ = {’cz’, ’id’, ’rx’, ’rz’, ’rzz’, ’sx’, ’x’}, which corresponds to the native gate set of the IBM quantum processing unit (QPU) ’ibm_fez’.
@@
We observed that strategically balancing the GHZ-task dataset with respect to both fidelity distribution and circuit length yielded substantial improvements in model performance. Specifically, the dataset was constructed with 25% of circuits having fidelity below 0.9, balanced evenly across different circuit lengths, while the remaining 85% consisted of circuits with high fidelity ($>0.9$) included without additional length balancing. Furthermore, we restricted the number of gates per circuit, using 3–16 gates for $gs1$ and 3–24 gates for $gs2$ and the  ML optimization task. Empirically, this targeted balancing approach notably improved training stability and enhanced the quality of generated quantum circuits.
@@
For the machine learning setting, SQuASH provides a dedicated set of quantum circuits along with their performance on the LinearlySeparable dataset introduced in [34], which is a linear classification task. As in the previous task, we define a specific gate set, {’cx’, ’h’, ’rx’, ’ry’, ’swap’, ’crx’, ’cry’} and strategically balance the dataset in terms of classification accuracy and circuit length, which significantly improves training performance. Similar to the quantum state preparation, we constructed the dataset so that 25% of circuits had accuracy below 0.75, evenly balanced across various circuit lengths, while the remaining 75% consisted of high-accuracy circuits ($>0.75$) included without further length balancing. Empirical results confirm that this targeted dataset composition notably enhances dataset quality,
IV-B GHZ Task:

### v3

@@ I Introduction
The promise of quantum computing hinges on the efficient design of quantum circuits executable on noisy intermediate-scale quantum (NISQ) hardware [2]. This is challenging due to hardware constraints like limited qubit counts and restricted gate sets [3, 4, 5]. The design of an appropriate ansatz for a given task – be it approximating a target unitary, solving an optimization problem or learning from data for a quantum machine learning (QML) task – is challenging and typically relies on human expertise or brute-force search. In recent years, there has been growing interest in leveraging heuristic optimization and machine learning (ML) techniques to automate quantum circuit design. Prior approaches include evolutionary strategies [6, 7, 8, 9], reinforcement learning [10, 11, 4, 12], Bayesian optimization [13], and adaptive algorithms [14, 15]. While these methods can discover novel circuit designs, they typically require repeated evaluations of quantum metrics or cost functions during training [16][17]. This evaluation process becomes intractable for large systems and trainable parameterized quantum circuits (PQCs), creating a significant computational bottleneck for automated circuit synthesis methods [18]. Moreover, techniques trained on specific quantum tasks or circuit layouts often show poor generalization to new scenarios, as they tend to encode task-specific features rather than learning transferable principles [19]. To address these challenges, [1] introduces the idea of applying denoising diffusion models (DM) to quantum circuit synthesis. The proposed text-conditioned diffusion pipeline generates gate-based circuits for a given target objective without explicitly simulating quantum dynamics during training. While this work established DMs as a promising tool for quantum circuit synthesis, an important open challenge is handling continuous gate parameters. Commonly used rotation gates, such as single-qubit or controlled rotations, are parameterized by continuous values and significantly impact the functionality of PQCs.
In this paper, we introduce an extended DM-based pipeline that integrates continuous-valued gate parameters alongside discrete gate choices into the generation process. We train the proposed model architecture on extensive datasets of quantum circuits with varying gate sets, designed for state preparation and QML classification tasks. Experimental results demonstrate that the described methodology effectively bypasses the computationally expensive parameter training phase during the search, addressing a key limitation of many existing approaches to the automated design of PQCs, and shows a strong ability to generalize to new scenarios. At inference time, the model, guided by a textual prompt, produces PQCs that are immediately deployable on the target quantum device, subject to specified constraints.
@@ II Related Work
The approach introduced in this work extends [1], which treats quantum circuits as structured data samples and trains a generative denoising DM to generate new circuit designs. DMs generate synthetic data by iteratively denoising random noise samples through a learned reverse diffusion procedure. Introduced by [20], the fundamental principle involves adding Gaussian noise to training samples as a forward Markovian diffusion process, then training a neural network to reverse the noising procedure and approximate the data distribution’s score function. After training, data generation begins from a pure noise sample, and the trained model iteratively predicts and removes noise, step by step, resulting in samples resembling those from the training dataset. DMs employ denoising autoencoder architectures implemented through neural networks, typically structured to predict either the original noise-free data or the noise itself. Recent advances, such as latent diffusion [21] and classifier-free guidance (CFG), have made DMs a leading approach in conditional generative modeling tasks across numerous domains, including image and video generation [21, 22], audio synthesis [23], and protein structure prediction [24]. Extending these advances, [1] explored the potential of DMs beyond classical domains and demonstrated their effectiveness in the context of quantum circuit synthesis. Specifically, the authors applied a denoising DM to two quantum computing tasks: (1) entanglement generation, where the model produced circuits that realize specific multi-qubit entangled states, and (2) unitary compilation, where the model generated circuits approximating random target unitaries using constrained gate sets. The DM excelled in generating valid, novel circuits and enabled advanced capabilities such as masking and editing parts of a circuit post hoc without retraining. Notably, the approach eliminates the need for classical simulations to compute cost functions comparing generated circuits to target unitaries. Ref. [25] proposes an architectural enhancement by replacing the standard U-Net used for noise prediction with a U-Net-style Diffusion Transformer. A recent work [26] applied a DM to generate ansätze for Variational Quantum Eigensolver (VQE), with a focus on chemistry problems. The model was trained on a corpus of unitary coupled cluster (UCC) ansatz circuits encoded into images and successfully proposed new circuits that achieved comparable ground-state energies. However, the techniques proposed in earlier studies represent gates using fixed-dimensional vectors drawn from a finite embedding set [27], limiting its ability to express parametrized quantum operations. In this work, we extend the methodology to enable the generation of PQCs with continuous gate parameters.
@@
Beyond diffusion-based models, this work also relates more broadly to the field of quantum architecture search (QAS) research, where the goal is to automatically generate PQCs tailored to specific tasks and hardware constraints [28]. While the most approaches proposed for QAS treat the design of PQCs as a two-step process [19, 4, 17, 13], i.e., first selecting discrete gate sequences and then using classical optimizers to determine continuous parameters, several recent methods aim to learn both circuit structures and parameterizations simultaneously, optimizing over both discrete gates and continuous parameters in a unified pipeline. Ref. [29] incorporates the parameter optimization task into reinforcement learning settings, enabling continuous actions. The generative quantum eigensolver (GQE) proposed in [30] is a transformer-based approach for generating PQCs. This approach is designed for the chemical problem of searching for the ground state of the electronic structure Hamiltonians of several molecules. Similarly to our approach, it integrates continuous optimizable gate parameters into a classical generative model. However, unlike the transformer-based model in GQE, we employ a DM for PQCs generation, which iteratively refines random noise samples to generate circuits, offering advantages in generating diverse quantum circuits with novel structures.
@@ III Methods
Figure 1: Overview of our diffusion-based PQC synthesis pipeline, inspired by [1] (a) Circuit Encoding: A circuit is converted into a tensor by concatenating the embeddings of discrete gates and continuous parameters (e.g., rotation angles). (b) Forward Diffusion: Gaussian noise is progressively injected into the tensor according to a predefined scheduler, emulating a corruption process. (c) Textual Conditioning: A pretrained CLIP encoder converts task-specific prompts into embeddings. (d) Reverse Diffusion: A U-Net with cross-attention receives the noisy tensor and text embedding, iteratively removing noise to restore gate identities and parameters. (e) Inference and Decoding: At inference, random noise is denoised under the given prompt, and the final tensor is mapped back into a readable quantum circuit.
@@ III-A Circuit Encoding
To train our denoising DM, we represent each PQC using two structured tensors, as depicted in Fig. 1 (a). The first tensor is three-dimensional, where one dimension corresponds to qubit indices $N$, another to discrete time steps $T$, and the third dimension encodes discrete gate identities $d_{c}$. For each discrete time step $t_{i}\in T$, only one gate per qubit is allowed. If no gate acts on a qubit at a particular timestep (i.e., an idle operation), we assign a special no-op token. Multi–qubit gates are represented by role–specific tokens placed on all participating qubits in the same time slot (e.g., a ’cx’ acting on qubits $i$ and $j$, where ’-’ (minus) represents target and ’+’ (plus) represents control). To convert this discrete representation into a continuous embedding suitable for the DM, each discrete gate type is transformed into a fixed-length embedding vector determined before training. This embedding process results in a continuous tensor representation of gates across qubits and time steps. To handle gates with continuous parameters, e.g., a rotation gate ’rx($\theta$)’, we introduce a second tensor with dimensions corresponding to qubit indices, discrete time steps, and the maximal number of normalized parameters used by any gate in the circuit $d_{p}$. Note that currently, our approach supports parameter embeddings with $d_{p}\leq 1$. In this parameter tensor, we encode the numerical values of gate parameters, placing a no-op token for gates without parameters. By concatenating these two tensors, each gate operation for every qubit and timestep is uniquely encoded by its discrete type embedding and its continuous parameter. As a result, identical gates with identical parameters map to identical points in this combined embedding space, whereas gates differing in type or parameter values map distinctly. This method naturally extends to gates with multiple parameters. During decoding, the combined tensor is split into a circuit matrix and a parameter matrix. For each qubit–timestep position in the circuit matrix, we use cosine similarity to match the generated gate embedding to its closest counterpart among the known gate embeddings. The corresponding parameter matrix is then applied only when a gate requires parameters, using the remaining $d_{p}$ dimensions of the vector as a parameter encoding. In principle, if a gate supports multiple parameters (not present in our current gate sets), one could introduce an inverse mapping or train an auxiliary decoder network to handle this more complex parameter space. If the chosen gate is non-parameterized (e.g., an ’h’ gate), the parameter subvector is simply ignored.
@@ III-B Model Architecture
In this work, we extend the quantum circuit synthesis pipeline described by [1]. Specifically, we introduce textual conditioning via a pre-trained  Contrastive Language-Image Pre-training (CLIP) text encoder - originally trained to align visual and textual representations - to guide the DM towards generating PQCs optimized for specific tasks, i.e., quantum state fidelity or classification accuracy maximization (Fig. 1 (c)). Task-specific textual prompts, for instance "Generate GHZ fidelity: $\mathsf{fid\_value}$” or "Generate Accuracy: $\mathsf{acc\_value}$”, are encoded into continuous embeddings to condition the diffusion process.
The core of our diffusion model utilizes a U-Net architecture, widely adopted in state-of-the-art diffusion-based image synthesis. The model accepts as input a noisy quantum circuit tensor of dimensions $(d_{c}+d_{p})\times N\times T$, representing embedding dimensions for gates and parameters, qubits and time steps, respectively. It outputs a noise-residual tensor of matching shape. To enable textual conditioning, we augment the standard U-Net structure with cross-attention layers, integrating CLIP-generated textual embeddings. This design explicitly conditions the generated circuits on task-specific textual prompts, eliminating the need for explicit labels of fidelity or accuracy, which the model implicitly learns from training data alone.
@@ III-C Training Strategy
Our DM is trained by simulating a forward diffusion process on quantum circuit data (Fig. 1 (b)) and learning its reverse denoising counterpart in a manner consistent with Denoising Diffusion Probabilistic Model (DDPM) [31] procedures (Fig. 1 (d)). Following the approach in [1], we parameterize the model as an $\epsilon$-predictor: at each diffusion timestep $t$, the model learns to directly predict the noise $\epsilon_{t}$ in a noisy tensor $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon_{t}\sim\mathcal{N}(0,I)$ (i.e., $\epsilon_{t}$ is drawn from a Gaussian distribution with zero mean and identity covariance), $x_{0}$ is a sample from the training dataset with arbitrary distribution $q(x_{0})$, and $\bar{\alpha}_{t}$ is the cumulative product of the variance schedule (i.e., $\bar{\alpha}_{t}=\prod_{i=0}^{t}(1-\beta_{i})$). The U-Net predicts this noise as $\hat{\epsilon}_{\theta}(x_{t},t,\texttt{prompt})$, conditioned on the task prompt. Model parameters are optimized to minimize the mean-squared error loss:
$\mathcal{L}=\mathbb{E}_{x_{0},\epsilon,t}\left[\left\|\epsilon-\hat{\epsilon}_{\theta}(x_{t},t,\texttt{prompt})\right\|^{2}\right],$
ensuring the network learns to denoise samples at all diffusion steps [1].
@@
Conditioning on textual prompts (e.g., ”Generate GHZ fidelity: 1.0000”) is done through cross-attention layers, where the prompt is embedded by a  CLIP encoder (output size $77\times 512$). For numerical stability, fidelity prompts are rounded to four decimal places, while QML accuracy prompts use one decimal place. We adopt this approach because small changes in prompt phrasing can induce noticeable variations in the generated circuits; thus, prompt sensitivity is mitigated by discretizing the input conditions. During training, with a 10% probability, the prompt is replaced by an empty token. This “unconditional” path enables CFG at inference, whereby we can smoothly interpolate between conditional and unconditional predictions [32]. We train for $T=1000$ diffusion steps using Adam optimizer with a one-cycle learning-rate schedule.
At inference time, we draw Gaussian noise as an initial circuit encoding and iteratively denoise it using the trained U-Net (Fig. 1 (e)). When CFG is enabled, we selectively mix conditional and unconditional predictions to control how strictly the model follows the prompt. The final denoised tensor is decoded into a human-readable quantum circuit: we pick the closest discrete gate embedding at each position and apply the predicted continuous parameter (if required) to obtain the final PQC.
@@ IV-A Dataset
A key element of our approach is the construction of large training datasets consisting of quantum circuits, which the diffusion model leverages to learn the circuit performance distribution. To address our two evaluation tasks, namely quantum state preparation and classification accuracy optimization, we utilize the existing SQuASH [^1] Available at https://github.com/SQuASH-bench/SQuASH dataset [33]. Specifically, for the quantum state preparation, the dataset includes quantum circuits designed to generate a $3$-qubit GHZ state within a predefined approximation error. The data were collected from various search strategies used to generate circuits for the task, including random search, reinforcement learning, and an evolutionary approach. To demonstrate the model’s adaptability to different hardware constraints, we define two distinct data subsets, each characterized by a different gate alphabet. The first subset incorporates PQCs constructed using the gate set $gs1$ = {’cx’, ’h’, ’rx’, ’ry’, ’rz’, ’id’}. The second subset uses $gs2$ = {’cz’, ’id’, ’rx’, ’rz’, ’rzz’, ’sx’, ’x’}, which corresponds to the native gate set of the IBM quantum processing unit (QPU) ’ibm_fez’.
@@
We observed that strategically balancing the GHZ-task datasets with respect to both fidelity distribution and circuit length yielded substantial improvements in model performance. Specifically, in $gs1$, 75% of the circuits had high fidelity ($>0.9$) without additional length balancing, while the remaining 25% had low fidelity ($\leq 0.9$) evenly distributed across all circuit lengths. In $gs2$, 90% of the circuits had high fidelity without length balancing, and the other 10% had low fidelity evenly distributed across the full range of circuit lengths. Furthermore, we restricted the number of gates per circuit to 3–16 for $gs1$ and 3–24 for $gs2$. Empirically, this targeted balancing approach notably improved training stability and enhanced the quality of generated quantum circuits.
@@
For the machine learning setting, SQuASH provides a dedicated set of quantum circuits along with their performance on the LinearlySeparable 222 https://github.com/XanaduAI/qml-benchmarks/blob/main/src/qml_benchmarks/data/linearly_separable.py dataset introduced in [34], which is a linear classification task. The dataset consists of 300 samples, each with $d=8$ features, and a predefined margin of $m=1$ between the two classes. To ensure reproducibility, we fix the random seed to 274 when generating the data. As in the previous task, we define a specific gate set, {’cx’, ’h’, ’rx’, ’ry’, ’swap’, ’crx’, ’cry’}, and strategically balance the dataset in terms of classification accuracy and circuit length, which significantly improves training performance. Similar to the quantum state preparation, we constructed the dataset so that 50% of circuits had accuracy below 0.7, evenly balanced across various circuit lengths, while the remaining 50% consisted of high-accuracy circuits ($\geq 0.7$) without further length balancing. All circuits were constrained to 3-24 gates. Empirical results confirm that this targeted dataset composition notably enhances model performance.
IV-B GHZ Task